OpenStudy (anonymous):

given log2=p and log7=r express log 14 in terms of p and r.

OpenStudy (anonymous):

log14=log2 + log 7 = p +r

OpenStudy (anonymous):

log(2*7) = log(2)+log(7) = p+r

OpenStudy (anonymous):

log14=log(2*7)=log2+log7=p+r

OpenStudy (anonymous):

thank you! <3